OpenStudy (jamanorthi):

solve 12(6-3.5)power 2

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=12%286-3.5%29%5E2%3D Heres a great website than can check answers, but your problem is solved my following order of operations.

OpenStudy (anonymous):

So for steps, first solve within parenthesis (6-3.5) = (2.5) (2.5)^2 = (6.25) 12(6.25) = 75 :) Any q's?

OpenStudy (anonymous):

@Jamanorthi ???